OpenStudy (anonymous):

this time i don't know the answer, but i give a medal. What is the diameter of a cone with height 8 m and volume 150m3 ? (1 point) 7.5 m 15 m

OpenStudy (anonymous):

A: 7.5 m B: 5 sqrt 3 m C: 7.5 sqrt 3 m d: 15 m

OpenStudy (anonymous):

V=(1/3)(pi)(r^2)(h) 150=(1.05)(r^2)(8) 150=(8.38)(r^2) 150/8.38=r^2 17.9=r^2 sqrt(17.9)=r 4.23=r d=2r d=8.46

OpenStudy (anonymous):

?

OpenStudy (anonymous):

volume of cone = 1/3 Pi r^2 h so: 150=1/3* Pi *r^2* 8 so r^2=450/(8Pi)=18 r=sqrt18 and diameter = 2sqrt18=6sqrt2

OpenStudy (dayianaforever):

answer 15